Find the equation of the line through the given pair of points. solve for y if possible. (2,-6) (7,2) I have to find the equation of the line and write it in standard form.
Answer by rothauserc(4718)

slope of line is the difference in y values divided by difference in x values,
we are given two points (2,-6) (7,2), therefore
slope = (2 - (-6)) / (7 - 2) = 8/5 = 1.6
*****************************************************************************
standard form for equation of a line is, y = mx + b where m is slope and b is the y intercept
so far we have
y = 1.6x + b
we can choose either point to determine b, let's choose (7,2)
2 = (1.6*7) + b
2 = 11.2 + b
subtract 11.2 from both sides of =
b = -9.2
therefore the standard form of the equation of the line is
y = 1.6x - 9.2